Circular No. 6348

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GRS 1739–278 J. Paul, L. Bouchet, E. Churazov, and R. Sunyaev, on behalf of the

SIGMA/Granat team (Service d’Astrophysique du CEA, Saclay; Centre d’Etude Spatiale des Rayonnements, Toulouse; Space Research Institute, Moscow), communicate: “During an observation of the galactic-center region started on Mar. 18 (effective duration 86 000 s), the SIGMA telescope onboard Granat detected a new hard x-ray source at the most probable position α = 17^{h}39^{m}28^{s}, δ = −27^{o}45′.6 (equinox 1950.0; uncertainty ± 3′). Because of some ambiguity inherent to coded-mask telescope imaging, the source position could also be at α = 17^{h}38^{m}07^{s}, δ = −34^{o}24′.0 (± 3′). Preliminary estimates of the flux level give 80 ± 28 mCrab in the range 40–75 keV, 51 ± 20 mCrab over 75–150 keV, and 104 ± 50 mCrab over 150–300 keV. A 5.5σ excess was detected at the same positions in the lower energy band during the previous observation (started on Mar. 16.77 UT), showing that the source was already active and probably brighter. Because of the hard spectrum, the source could be a new blackhole x-ray nova, and observations at other frequencies are urged. Further SIGMA observations should clarify position ambiguity.”

PKS 0208–512 W. T. Vestrand and J. G. Stacy, University of New Hampshire; and

R. Mukherjee and P. Sreekumar, Goddard Space Flight Center, NASA, report: “Quick-look analysis of data taken during Mar. 5–16 by the Energetic Gamma Ray Experiment Telescope (EGRET) aboard the Compton Gamma Ray Observatory indicates that PKS 0208–512 is currently in a high-intensity state. The current integral flux above 100 MeV is > 10^{−6} photon cm^{−2} s^{−1} and is comparable to or exceeds the highest γ-ray flux ever seen from this variable γ-ray blazar. While quick-look data are not yet available for flux estimates at MeV energies, PKS 0208–512 has been proposed as a prototype for a class of ‘MeV blazars’ that exhibit highly variable emission at MeV energies. Multi-wavelength observations are essential for modeling this potential new class of objects, and we encourage southernhemisphere observers to observe PKS 0208–512 while in outburst.”

COMET C/1995 O1 (HALE-BOPP) Visual m_{1} estimates by R. Keen, Mt. Thorodin, CO: Feb. 24.52 UT,

8.6 (0.15-m reflector); Mar. 15.50, 8.5; 20.50, 8.6 (0.32-m reflector).

1996 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green